OpenStudy (anonymous):

When n = energy efficiency and Pin = energy input and Pout = energy output, how can you mathematically represent the correct relationship between the energy you put in something and the energy you get out? A. n = P out * P in B. n = P out / P in C. n = P out - P in

OpenStudy (anonymous):

efficiency is equal to the ratio of energy in and out. efficiency will be equal to or less than one. does this help?

OpenStudy (anonymous):

yes and no which one is the correct answer??

OpenStudy (anonymous):

openstudy doesn't work that way. I'll teach you if you want to learn how to know which is the correct answer, if you want.

OpenStudy (anonymous):

if you have the time to help then i would appreciate that

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

efficiency is a ratio of how much energy is put in to how much energy is put out. for example if I asked you what percent is 90 or 100? you'd say..?

OpenStudy (anonymous):

90%? or 100%

OpenStudy (anonymous):

i dont really know ive never been good at math

OpenStudy (anonymous):

yaya, you got it!

OpenStudy (anonymous):

Sorry, i was helping another person ^_^

OpenStudy (anonymous):

right it would be 90% efficiency is just a percent. so to get 90% did you divide or times or subtract?

OpenStudy (anonymous):

90*100 = 9000 [so not times] 90/100 = 0.90 = 90% [so divide] 90-100 = -10 [so not subtraction]

OpenStudy (anonymous):

OK thank you i think i got the hang of this thanks for your help i appreciate it

OpenStudy (anonymous):

yep